1. Let g(t) = We are given g(t) 2-3 31(12-9) (12 - 3)2 and (t) 186(+9) (t2-3) (a) Determine the domain of g. (b) Determine
Answer #1

ght) = x² t²3 0 @ To find domain of get) Denominator #0 - 3 0 =) t+18) (tr) to w Domain - R-f-8.8 © =0 14) >0 ght 29 gire) =t: +3) git) = 3(427) 62 = 13.5 maximum = 13.5 © g(t) = 18 41449) (t2-33 points gullt)=0, To find inflection os 18+ (479) =0